# Low Complexity Channel Estimation for RIS-Assisted THz Systems with Beam Split

### v1

@@ IV-B DS-MUSIC Based Coupled DoAs/DoDs Estimation at RIS
$\tilde{p}_{k,\bar{m},l,q}^{\mathrm{MUSIC}}=\frac{1}{\left\|(\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}})^{\mathsf{H}}\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\right\|_{2}^{2}},q=1,\dots,Q_{\mathrm{B}},$ (69)
@@
where $\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\!\triangleq\!e^{\!-j2\pi\eta_{m}\left(\!\frac{q}{Q_{\mathrm{R}}}\!-1\right)\mathbf{n}_{\mathrm{sub}}}$ with $\mathbf{n}_{\mathrm{sub}}\!=\!\left[1,\cdots,N_{\mathrm{sub}}\right]$. The MUSIC-PS is obtained by mapping the sampled beams $\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}$ onto the noise space $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}$. Based on the fact that any beam $\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\in\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}$ that resides in the signal subspace is orthogonal to the noise subspace $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}$, the $\ell_{2}$-norm in (69) associated with the actual directions should approximate zero. The MUSIC algorithm is typically dependent on peak detection to find actual physical directions, resulting in undetectable marginal peaks. To address this issue, we design a novel extended-log-MUSIC-PS (EL-MUSIC-PS), where we extend the spectrum and map it to the logarithmic domain to enhance the accuracy of estimation. The EL-MUSIC-PS can be express as
$\mathbf{p}_{k,\bar{m},l}^{\mathrm{MUSIC}}=\log\left[\min\left(\tilde{\mathbf{p}}_{k,\bar{m},l}^{\mathrm{MUSIC}}\right),\tilde{\mathbf{p}}_{k,\bar{m},l}^{\mathrm{MUSIC}},\min\left(\tilde{\mathbf{p}}_{k,\bar{m},l}^{\mathrm{MUSIC}}\right)\right].$ (70)
In this form, the actual directions can be determined by employing local-maximum algorithms. Recalling that the sampled beams $\xi_{\mathrm{R},m,q}^{\mathrm{sub}}$ can be written as
@@
$\begin{split}\xi_{\mathrm{R},m,q}^{\mathrm{sub}}&=e^{-j2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1\right)\mathbf{n}_{\mathrm{sub}}}\\ &=e^{-j2\pi K_{q}\mathbf{n}_{\mathrm{sub}}}e^{-j2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1\right)\mathbf{n}_{\mathrm{sub}}}\\ &=e^{-j2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1+\frac{K_{q}}{\eta_{m}}\right)\mathbf{n}_{\mathrm{sub}}},\end{split}$ (71)
@@
where $K_{q}\in\mathbb{N},\forall q=1,\dots,Q_{\mathrm{B}}$. It is observed that when $q^{\prime}=q+K_{q}Q_{\mathrm{R}}/\eta_{m}$, $\xi_{\mathrm{R},m,q}^{\mathrm{sub}}$ will generate the identical beam. Moreover, the maximum number of identical beams $K_{q}^{\max}$ can be obtained by solving the integer equation $K_{q}^{\max}Q_{\mathrm{R}}/\eta_{m}=Q_{\mathrm{B}}$, which can be derived as
$K_{q}^{\max}=\lfloor\eta_{m}\left(2-1/Q_{\mathrm{R}}\right)\rfloor+1.$ (72)
@@
Therefore, more than $J_{k}$ peaks would be detected in the EL-MUSIC-PS in (70), making it challenging to distinguish the true and fake physical directions in one single SC. Thankfully, it is observed from (72) that the spatial directions associated with true physical directions keep frequency-independent. Inspired by this, we propose a novel double sensing (DS) algorithm to acquire the true physical direction. In this approach, we select two SCs, calculate and compare their EL-MUSIC-PSs on the same plot. The positions where peaks overlap indicate the true physical directions. However, due to the unavoidable noise and estimation errors in full CSI, the EL-MUSIC-PSs of different SCs may exist offsets. Therefore, in practical applications, we select the $J_{k}$ closest peaks to represent true physical directions, and we introduce the peak distance $d$ to measure the distance between two peaks, which is given by
@@
where $q_{m}$ and $q_{m^{\prime}}^{\prime}$ denote the indices of the peaks at two selected SCs, respectively. It can be observed from (72) that the selection of SC $m$ affects the number of identical beams. To avoid additional algorithm complexity, the selection of SCs should satisfy the constraint $K_{q,\max}\leq 2$. Employing the floor function inequality $\lfloor x\rfloor\leq x$ and substituting (5) into it, we can obtain
@@
Substituting (5) into (73), the peak distance $d$ can be derived as
$d\left(q_{m},q_{m^{\prime}}^{\prime}\right)=K_{q}Q_{\mathrm{R}}f_{c}\left|\frac{1}{f_{m^{\prime}}}-\frac{1}{f_{m}}\right|.$ (75)
To distinguish true physical directions effectively, the peak distance should be sufficiently large. According to (75), it is evident that the intervals between selected SCs should be large enough. Therefore, under the constraint in (74), we select the first and the $M/2$-th SCs to enhance the accuracy of estimation. Fig. 4 depicts an example of estimated EL-MUSIC-PS with the proposed DS approach. Assuming that the cascaded channel contains $J_{k}=3$ RIS-UE paths for a arbitrary UE with SNR $=20$ dB. It can be observed that the peaks overlap in actual physical directions, while that associated with fake spatial directions are totally separated. The estimated physical directions can be easily distinguished from the fake ones and accurately match the actual physical directions.
Figure 4: Normalized EL-MUSIC-PS via DS-MUSIC algorithm.
IV-C LS Estimation for Angle-Excluded Coefficients
With the estimated DoDs at BS and coupled DoAs/DoDs at RIS, (10) can be rewritten as
@@ IV-C LS Estimation for Angle-Excluded Coefficients
where $\tilde{\mathbf{x}}_{k,m}\!\triangleq\!\boldsymbol{\beta}_{k,m}\!\otimes\!\mathbf{\Upsilon}_{m}\!\in\mathbb{C}^{L^{2}\!J_{k}}$ denotes the angle-excluded vector, $\hat{\boldsymbol{\varphi}}\triangleq[\hat{\varphi}_{1},\dots,\hat{\varphi}_{L}]\in\mathbb{C}^{L}$ denotes the estimated DoDs at BS of all $L$ paths, and $\hat{\boldsymbol{\gamma}}_{k}\triangleq[\hat{\boldsymbol{\gamma}}_{k,1}^{\mathsf{T}},\dots,\hat{\boldsymbol{\gamma}}_{k,L}^{\mathsf{T}}]^{\mathsf{T}}\in\mathbb{C}^{LJ_{k}}$ denotes the estimated coupled directions at RIS. Substituting (76) into (19), the received signal $\mathbf{y}_{k,t}\left[m\right]$ can be rewritten as
@@
where $\tilde{\mathbf{\Psi}}_{m}\!\triangleq\overline{\mathbf{R}}_{m}\left(\mathbf{A}_{\mathrm{T},m}^{*}\left(\hat{\boldsymbol{\varphi}}\right)\!\otimes\!\mathbf{A}_{\mathrm{R},m}\left(\hat{\boldsymbol{\gamma}}_{k}\right)\right)\!\in\!\mathbb{C}^{PT\times L^{2}J_{k}}$. In this case, the angle-excluded coefficients $\tilde{\mathbf{x}}_{k,m}$ can be easily obtained employing LS solution
@@
Then, the full CSI can be calculated by
@@
IV-D Complexity Analysis
In this subsection, we analyze the computational complexity of the proposed channel estimation scheme. Specifically, the complexity of the proposed scheme mainly stems from four major steps: 1) Full CSI estimation via CBS-GAMP approach $\mathcal{O}\left(2PTQ_{\mathrm{T}}Q_{\mathrm{B}}\right)$, 2) DoDs acquisition at BS via EnM method $\mathcal{O}\left(Q_{\mathrm{T}}Q_{\mathrm{B}}\right)$, 3) Coupled directions acquisition at RIS via DS-MUSIC algorithm $\mathcal{O}\left(2N_{r}^{3}\right)$, 4) Angle-excluded coefficients estimation via LS solution $\mathcal{O}\left(\left(M-2\right)L^{6}J_{k}^{3}\right)$. The overall complexity of the proposed scheme is given by
$\mathcal{O}\left(2PTQ_{\mathrm{T}}Q_{\mathrm{B}}+2N_{r}^{3}+\left(M-2\right)L^{6}J_{k}^{3}\right)$ (80)
For comparison, we also discuss the complexity of the conventional OMP scheme [33], BSA-OMP scheme [14] and CBS-GAMP scheme. Since the above algorithms need to be executed at each SC, the total computational complexity thus is multiplied by $M$ compared to the original. The conventional OMP algorithm has the same complexity $\mathcal{O}\left(M{Q_{\mathrm{T}}}^{3}{Q_{\mathrm{R}}}^{6}\right)$ as BSA-OMP approach. The computational complexity of CBS-GAMP is $\mathcal{O}\left(MPTQ_{\mathrm{T}}Q_{\mathrm{B}}\right)$. Due to the sparse characteristics of THz channel, the number of propagation path satisfies $L,J_{k}\ll Q_{\mathrm{T}},Q_{\mathrm{R}}$. The number of SCs $M$ is also quite large in wideband systems. Therefore, it is clear that the computational complexity of the proposed scheme is much lower than other three approaches.
@@ V Numerical Results and Discussion
In this section, we present numerical results to validate the effectiveness of the proposed scheme. Without loss of generality, we consider a wideband RIS-assisted THz scenario with $f_{c}=100$ GHz, $B=15$ GHz and $M=128$. A BS employing a $N_{\mathrm{T}}$ = 16 ULA transmits signals to $K=8$ single antenna UEs, with the assistance of a RIS equipped with $N_{\mathrm{R}}=64$ passive reflecting elements. For both BS-RIS and RIS-UEs channels, there exit $L=J_{k}=3,\forall k$ propagation paths and the AoAs and AoDs are uniformly generated from the discretized grid within the range $\left[-\pi/2,\pi/2\right]$. The quantization level of CBS dictionaries is set as $Q_{\mathrm{T}}=32$ and $Q_{\mathrm{R}}=128$. As a metric for CSI estimation performance, we use the normalized mean square error (NMSE), which is given by
$\mathrm{NMSE}=\mathbb{E}\left\{\frac{\left\|\mathbf{h}_{k}^{\mathrm{cas}}[m]-\mathbf{\hat{h}}_{k}^{\mathrm{cas}}[m]\right\|_{2}^{2}}{\left\|\mathbf{h}_{k}^{\mathrm{cas}}[m]\right\|_{2}^{2}}\right\}$ (81)
Besides, we employ the correct probability as a metric for direction estimation performance, which is given by
$\mathrm{P}_{\mathrm{c}}=\mathrm{Pr}\left\{\left|\hat{\theta}-\theta\right|<\varepsilon_{\mathrm{th}}\right\}$ (82)
where $\varepsilon_{\mathrm{th}}\!=\!\delta/2$ is the threshold which is set as half the sample interval. For comparison, we also consider the conventional OMP scheme [33], BSA-OMP scheme [14], CBS-GAMP scheme, and oracle-LS scheme as benchmark schemes in the simulation. It is notice that the angular information is assumed perfectly known in oracle-LS scheme, which can be regarded as the performance upper bound of the proposed scheme.
@@
Fig. 5 depicts the correct probability performance of the DoDs at BS and coupled DoAs/DoDs at RIS versus SNR with $T=50$ and $P=30$. It is shown that both the two correct probabilities approach one as SNR increases. This indicates that the proposed scheme can ensure a certain level of reliability even with low SNR conditions and has the ability to obtain accurate physical directions with sufficiently well channel conditions. Besides, the correct probability of the coupled directions at RIS is lower than that of the DoDs at BS in all SNR regions. This is because the estimation of the former relies on the latter. Angular estimation errors at BS would propagate to RIS, leading to performance gap between the two probability curves.
@@
Fig. 6 illustrates the NMSE performance versus SNR with $T=50$ and $P=30$. Overall, we can observe that the proposed scheme outperforms conventional-OMP and BSA-OMP algorithm in all SNR regions. This is because the proposed scheme has the ability to exploit unknown channel information, and the proposed dictionaries can effectively cover the beam space under beam split effect. Besides, it can be observed that the performance of the proposed scheme is slightly inferior to that of CBS-GAMP algorithm when SNR is less than $10$ dB. This is because estimation errors generated by partial SCs propagate more significantly to the remaining SCs under poor channel conditions. However, the proposed scheme outperforms CBS-GAMP algorithm in high SNR regions, i.e., SNR is larger than $10$ dB. This is because angle decoupling provides more accurate angular estimation information. This implies that there exists a performance-complexity trade-off between these two algorithms. Moreover, it is obvious that the proposed scheme can approach NMSE performance of the oracle-LS scheme, verifying the effectiveness of the proposed scheme.
@@
Fig. 7 shows the NMSE performance versus the number of subframes $T$, with SNR $=20$ dB and $P=15$. As expected, with the increase in the number of subframes $T$, NMSE performance of all the algorithms improves. We can observe that the proposed scheme can maintain well performance even in the case with a short length of subframes, e.g., $T=10$, demonstrating the robustness of the proposed algorithm. Moreover, It can be observed that compared to the CBS-GAMP method, the proposed algorithm can reach similar performance with approximately $50\%$ subframes reduction. This indicates that the proposed scheme can efficiently reduce the pilot overhead for channel estimation.
@@
Fig. 8 provides the NMSE performance versus the number of time slots $P$, with SNR $=20$ dB and $T=30$. As expected, with the increase in the number of slots $P$, NMSE performance of all the algorithms improves. We can observe that the proposed scheme outperforms all other three algorithms and can achieve near-optimal NMSE performance of the ideal oracle-LS upper bound in all considered $P$. The performance gap between the proposed scheme and other existing schemes is quite great, demonstrating that the proposed scheme can effectively reduce the system complexity for channel estimation.
@@
Fig. 9 illustrates the NMSE performance versus the number of paths $L$ and $J_{k}$, with $T=50$ and $P=30$. We can observe that as the the number of paths increases, the NMSE performance of both the two schemes degrade. This is because of the increased estimation parameters and challenges in distinguishing directions between more multipaths. Furthermore, as SNR increases, the NMSE of CBS-GAMP algorithm converges to around $5\times 10^{-4}$ regardless of the number of paths, while the NMSE of the proposed scheme remain decreasing in all SNR regions. This indicates that the proposed scheme exhibits a superior performance.
@@
Fig. 10 shows the NMSE performance versus the number of reflecting elements of RIS $N_{\mathrm{R}}$, with $T=50$ and $P=30$. We can observe that as $N_{\mathrm{R}}$ decreases, the NMSE performance of the BSA-OMP algorithm degrade greatly. This is because the small-scale RIS cannot provide enough measurement data diversity to support OMP based algorithms. In contrast, the performance of the proposed and CBS-GAMP schemes slightly improve as $N_{\mathrm{R}}$ decreases. This indicates that these two schemes exhibit robust framework, which are less susceptible to the lack of data diversity.

### v2

@@ III-C2 Common Coupled Angle Estimation at RIS
$\tilde{p}_{k,\bar{m},l,q}=\frac{1}{\left\|(\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}})^{\mathsf{H}}\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\right\|_{2}^{2}},q=1,\dots,Q_{\mathrm{B}}.$ (35)
@@
Given that any beam $\xi_{\mathrm{R},\bar{m},q}^{\mathrm{sub}}\in\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{S}}$ in the signal subspace should ideally have minimal projection onto the noise subspace $\mathbf{\Omega}_{k,\bar{m},l}^{\mathrm{N}}$, which means that the $\ell_{2}$-norm in (35) associated with the actual directions should approximate zero. The MUSIC algorithm typically relies on peak detection to find actual physical directions. To avoid undetectable marginal peaks, we extend the pseudo-spectrum by inserting the minimum value at its edges. The extended pseudo-spectrum can be express as
$\mathbf{p}_{k,\bar{m},l}=\left[\min\left(\tilde{\mathbf{p}}_{k,\bar{m},l}\right),\tilde{\mathbf{p}}_{k,\bar{m},l},\min\left(\tilde{\mathbf{p}}_{k,\bar{m},l}\right)\right].$ (36)
The potential directions can then be determined by employing local-maximum algorithms. However, due to the extended range of coupled angles, the estimated number of directions will be $K_{\text{cp}}$ times the original.
Proposition 2.
The maximum number of identical beams can be given as
$K_{\rm{cp},\max}=\lfloor\eta_{m}\left(2-1/Q_{\mathrm{R}}\right)\rfloor+1.$ (37)
Proof.
Recalling that the sampled beams $\xi_{\mathrm{R},m,q}^{\mathrm{sub}}$ can be written as
@@
$\begin{split}\xi_{\mathrm{R},m,q}^{\mathrm{sub}}&=e^{-j2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1\right)\mathbf{n}_{\mathrm{sub}}}\\ &=e^{-j2\pi K_{\text{cp}}\mathbf{n}_{\mathrm{sub}}}e^{-j2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1\right)\mathbf{n}_{\mathrm{sub}}}\\ &=e^{-j2\pi\eta_{m}\left(\frac{q}{Q_{\mathrm{R}}}-1+\frac{K_{\text{cp}}}{\eta_{m}}\right)\mathbf{n}_{\mathrm{sub}}},\end{split}$ (38)
@@
where $K_{\text{cp}}\in\mathbb{N},\forall q=1,\dots,Q_{\mathrm{B}}$. It is observed that when $q^{\prime}=q+K_{\text{cp}}Q_{\mathrm{R}}/\eta_{m}$, the element in sub-CBS dictionary will generate the identical beam, i.e., $\xi_{\mathrm{R},m,q}^{\mathrm{sub}}=\xi_{\mathrm{R},m,q^{\prime}}^{\mathrm{sub}}$. Therefore, the maximum number of identical beams $K_{\text{cp},\max}$ can be obtained by solving the integer equation $K_{\text{cp}}Q_{\mathrm{R}}/\eta_{m}=Q_{\mathrm{B}}$. By rearranging the terms and employing the relationship $Q_{\mathrm{B}}=2Q_{\mathrm{R}}-1$, we can obtain
$K_{\text{cp}}=\frac{\left(2Q_{\mathrm{R}}-1\right)\eta_{m}}{Q_{\mathrm{R}}}.$ (39)
Therefore, the maximum value of $K_{\text{cp}}$ is the integer part of this expression, which can be given as
$K_{\rm{cp},\max}=\left\lfloor\eta_{m}\left(2-\frac{1}{Q_{\mathrm{R}}}\right)\right\rfloor+1.$ (40)
@@
According to Proposition 2, instead of the expected $J_{k}$ peaks in the MUSIC pseudo-spectrum, $K_{\text{cp},\max}J_{k}$ peaks will be detected. This makes the algorithm infeasible for distinguishing the true physical directions from fake ones within a single SC. Thankfully, it is observed from (37) that the spatial directions associated with true physical directions keep frequency-independent. Inspired by this, we propose a novel double sensing (DS) algorithm to acquire the true physical direction. In this approach, we select two SCs to calculate and compare their extended MUSIC pseudo-spectrum. The positions where peaks overlap indicate the true physical directions. However, due to the unavoidable noise and estimation errors, the pseudo-spectrum of different SCs may exist offsets. Therefore, in practical implement, we select the $J_{k}$ closest peaks to represent true physical directions. We introduce the peak distance $d$ to measure the distance between two peaks, which is given by
@@
where $q_{m}$ and $q_{m^{\prime}}^{\prime}$ denote the indices of the peaks at two selected SCs, respectively. So far, the detailed steps of the proposed DS-MUSIC algorithm are summarized in Algorithm 2.
To ensure the algorithm is both effective and efficient, it is crucial to select the appropriate SCs. Since $K_{\text{cp},\max}$ in (37) is affected by the choice of SCs, we should aim to satisfy $K_{\text{cp},\max}\leq 2$, which means each angle has at most two identical peaks. Employing the floor function inequality $\lfloor x\rfloor\leq x$ and substituting (5) into (37), we can obtain
@@
Besides, substituting (5) into (41), the peak distance $d$ can be derived as
$d\left(q_{m},q_{m^{\prime}}^{\prime}\right)=K_{\text{cp}}Q_{\mathrm{R}}f_{c}\left|\frac{1}{f_{m^{\prime}}}-\frac{1}{f_{m}}\right|.$ (43)
To distinguish true physical directions effectively, the peak distance should also be sufficiently large. According to (43), it is evident that the intervals between selected SCs should be large enough. Therefore, under the constraint in (42), we select the first and the $M/2$-th SCs to enhance the accuracy of estimation.
Fig. 5 depicts an example of estimated extended MUSIC pseudo-spectrum with the proposed DS approach. Assuming that the cascaded channel contains $J_{k}=3$ RIS-UE paths for an arbitrary UE with SNR $=20$ dB. It can be observed that the peaks overlap at actual physical directions, while those associated with fake spatial directions are completely separated. The estimated physical directions can be easily distinguished from the fake ones and accurately match the actual physical directions.
Figure 5: Normalized extended MUSIC pseudo-spectrum via DS-MUSIC algorithm.
III-D Channel Estimation for Other Subcarriers
Through the proposed two-step full CSI estimation, we obtained the cascaded channels for two subcarriers and the corresponding common physical angle. Based on this knowledge, the cascaded channel in (10) can be rewritten as
@@ III-D Channel Estimation for Other Subcarriers
where $\tilde{\mathbf{x}}_{k,m}\!\triangleq\!\boldsymbol{\beta}_{k,m}\!\otimes\!\mathbf{\Sigma}_{m}\!\in\mathbb{C}^{L^{2}\!J_{k}}$ represents the angle-excluded vector, $\hat{\boldsymbol{\varphi}}\triangleq[\hat{\varphi}_{1},\dots,\hat{\varphi}_{L}]\in\mathbb{C}^{L}$ is the estimated DoDs at BS of all $L$ paths, and $\hat{\boldsymbol{\gamma}}_{k}\triangleq[\hat{\boldsymbol{\gamma}}_{k,1}^{\mathsf{T}},\dots,\hat{\boldsymbol{\gamma}}_{k,L}^{\mathsf{T}}]^{\mathsf{T}}\in\mathbb{C}^{LJ_{k}}$ denotes the estimated coupled directions at RIS.
Substituting (44) into (20), the received signal $\mathbf{y}_{k,t}\left[m\right]$ can be rewritten as
@@
where $\tilde{\mathbf{\Psi}}_{m}\!\triangleq\bar{\mathbf{R}}_{m}\left(\mathbf{A}_{\mathrm{T},m}^{*}\left(\hat{\boldsymbol{\varphi}}\right)\!\otimes\!\mathbf{A}_{\mathrm{R},m}\left(\hat{\boldsymbol{\gamma}}_{k}\right)\right)\!\in\!\mathbb{C}^{PT\times L^{2}J_{k}}$. The column size of the $\tilde{\mathbf{\Psi}}_{m}$ matrix is reduce from $Q_{\mathrm{B}}Q_{\mathrm{T}}$ of overcomplete dictionary to a multipath-related size $L^{2}J_{k}$. This significant reduction in the column size enhances both the efficiency and feasibility of the LS solution. Therefore, the angle-excluded coefficients $\tilde{\mathbf{x}}_{k,m}$ can be easily obtained as
@@
Substituting (46) into (44), the full CSI for other SCs $m\in\mathcal{M}\backslash\{1,M/2\}$ can be expressed as
@@
IV Analysis of Computational Complexity
In this section, we analyze the computational complexity of the proposed channel estimation protocol. We also compare our results with the other existing algorithms. Specifically, the complexity of the proposed scheme mainly stems from four major components:
  1. 1.
    Cascaded channel estimation for two selected SCs by CBS-based GAMP approach: $\mathcal{O}\left(2PTQ_{\mathrm{T}}Q_{\mathrm{B}}\right)$.
  2. 2.
    DoDs acquisition at BS: $\mathcal{O}\left(Q_{\mathrm{T}}Q_{\mathrm{B}}\right)$.
  3. 3.
    Coupled directions acquisition at the RIS using DS-MUSIC algorithm: $\mathcal{O}\left(2N_{r}^{3}\right)$.
  4. 4.
    Channel estimation for other SCs: $\mathcal{O}\left(\left(M-2\right)L^{6}J_{k}^{3}\right)$.
The overall complexity of the proposed scheme can be expressed as
$\mathcal{O}\left(2PTQ_{\mathrm{T}}Q_{\mathrm{B}}+2N_{r}^{3}+\left(M-2\right)L^{6}J_{k}^{3}\right).$ (48)
For comparison, we also discuss the complexity of the conventional OMP scheme [30], BSA-OMP scheme [17] and CBS-GAMP scheme [23]. These algorithms perform under traditional channel estimation frameworks which means they need to be executed at each SC or simultaneously process all SCs. The computational complexity of conventional OMP algorithm is $\mathcal{O}\left(M{Q_{\mathrm{T}}}^{3}{Q_{\mathrm{R}}}^{6}\right)$. The BSA-OMP approach has the same complexity as the conventional OMP, $\mathcal{O}\left(M{Q_{\mathrm{T}}}^{3}{Q_{\mathrm{R}}}^{6}\right)$, since it similarly processes each subcarrier independently. The computational complexity of CBS-GAMP is $\mathcal{O}\left(MPTQ_{\mathrm{T}}Q_{\mathrm{B}}\right)$. Additionally, the number of SCs $M$ is typically large in wideband systems. Therefore, it is evident that the computational complexity of the proposed scheme is significantly lower than the other approaches mentioned.
@@ V Numerical Results and Discussion
In this section, we present numerical results to validate the effectiveness of the proposed scheme. Without loss of generality, we consider a wideband RIS-assisted THz scenario with central carrier frequency $f_{c}=100$ GHz, bandwidth $B=15$ GHz and $M=128$ SCs. A BS employing a ULA with $N_{\mathrm{T}}$ = 16 antennas transmits signals to $K=8$ single antenna UEs, with the assistance of a RIS equipped with $N_{\mathrm{R}}=64$ passive reflecting elements. For both BS-RIS and RIS-UEs channels, there exit $L=J_{k}=3,\forall k$ propagation paths, and the AoAs and AoDs are randomly generated from the discretized grid within the range $\left[-\pi/2,\pi/2\right]$. The quantization level of overcomplete dictionaries is set as $Q_{\mathrm{T}}=32$ and $Q_{\mathrm{R}}=128$.
We use the normalized mean square error (NMSE) as a metric for CSI estimation performance, which is given by
$\mathrm{NMSE}=\mathbb{E}\left\{\frac{\left\|\mathbf{h}_{k}^{\mathrm{cas}}[m]-\mathbf{\hat{h}}_{k}^{\mathrm{cas}}[m]\right\|_{2}^{2}}{\left\|\mathbf{h}_{k}^{\mathrm{cas}}[m]\right\|_{2}^{2}}\right\}.$ (49)
Additionally, we employ the correct probability as a metric for direction estimation performance, which is given by
$\mathrm{P}_{\mathrm{c}}=\mathrm{Pr}\left\{\left|\hat{\theta}-\theta\right|<\varepsilon_{\mathrm{th}}\right\},$ (50)
where $\varepsilon_{\mathrm{th}}\!=\!\delta/2$ is the threshold which is set as half the sample interval. For comparison, we also consider the conventional OMP scheme [30], BSA-OMP scheme [17], CBS-GAMP scheme [23], and oracle-LS scheme [31] as benchmark schemes in the simulation. It is noted that the angular information is assumed perfectly known in oracle-LS scheme, which can be regarded as the performance upper bound of the proposed scheme.
@@
Fig. 6 depicts the correct probability performance of the DoDs at BS and coupled DoAs/DoDs at RIS versus SNR with $T=50$ and $P=30$. It is shown that both the correct probabilities approach one as SNR increases. This indicates that the proposed scheme can ensure a certain level of reliability even under low SNR conditions and has the ability to obtain accurate physical directions with sufficient channel conditions. Additionally, the correct probability of the coupled directions at RIS is consistently lower than that of the DoDs at BS across all SNR values. This is because the estimation of the coupled directions at RIS depends on the accuracy of the DoDs estimation at BS. Angular estimation errors at BS would propagate to RIS, leading to performance gap between the two probability curves.
@@
Fig. 7 illustrates the NMSE performance versus SNR with $T=50$ and $P=30$. Overall, it can be observed that the proposed scheme outperforms conventional OMP and BSA-OMP algorithm across all SNR regions. This is because the proposed scheme has the ability to exploit unknown channel information, and the proposed dictionaries can effectively cover the beam space under beam split effect. Additionally, the proposed scheme is slightly inferior to that of CBS-GAMP algorithm when SNR is less than $10$ dB. This is because estimation errors generated by partial SCs propagate more significantly to the remaining SCs under poor channel conditions. However, the proposed scheme outperforms CBS-GAMP algorithm in high SNR regions, i.e., SNR is larger than $10$ dB. This is because angle decoupling provides more accurate angular estimation information. This suggests that there exists a performance-complexity trade-off between these two algorithms. Moreover, it is obvious that the proposed scheme can approach the NMSE performance of oracle-LS scheme, which verifies the effectiveness of the proposed scheme.
@@
Fig. 8 shows the NMSE performance versus the number of subframes $T$, with SNR $=20$ dB and $P=15$. As expected, the NMSE performance of all the algorithms improves with the increase in the number of subframes $T$. It can be observed that the proposed scheme maintains good performance even with a short length of subframes, e.g., $T=10$, demonstrating the robustness of the proposed algorithm. Moreover, compared to the CBS-GAMP method, the proposed algorithm achieves similar performance with approximately a $50\%$ subframes reduction. This indicates that the proposed scheme can efficiently reduce the pilot overhead for channel estimation.
@@
Fig. 9 provides the NMSE performance versus the number of time slots $P$, with SNR $=20$ dB and $T=30$. As expected, the NMSE performance of all the algorithms improves with the increase in the number of slots $P$. It can be observed that the proposed scheme outperforms all other three algorithms and can achieve near-optimal NMSE performance of the ideal oracle-LS upper bound in all considered values of $P$. The performance gap between the proposed scheme and other existing schemes is quite large. This indicates that the proposed scheme can effectively reduce the system complexity for channel estimation.
@@
Fig. 10 illustrates the NMSE performance versus the number of paths $L$ and $J_{k}$, with $T=50$ and $P=30$. It can be observed that as the the number of paths increases, the NMSE performance of both the two schemes degrade. This is because of the increased estimation parameters and challenges in distinguishing directions among more multipath components. Furthermore, as SNR increases, the NMSE of CBS-GAMP algorithm converges to around $5\times 10^{-4}$ regardless of the number of paths. In contrast, the NMSE of the proposed scheme remain decreasing across all SNR regions. This demonstrates that the proposed scheme maintains better estimation accuracy and is more robust against variations in the number of paths compared to the CBS-GAMP algorithm.
@@
Fig. 11 shows the NMSE performance versus the number of reflecting elements of RIS $N_{\mathrm{R}}$, with $T=50$ and $P=30$. It can be observed that as $N_{\mathrm{R}}$ decreases, the NMSE performance of the BSA-OMP algorithm degrades significantly. This is because a small-scale RIS cannot provide enough measurement data diversity to support OMP based algorithms. In contrast, the performance of the proposed and CBS-GAMP schemes slightly improve as $N_{\mathrm{R}}$ decreases. This indicates that these two schemes exhibit robust framework, which are less susceptible to the lack of data diversity.